# Nonlinear Online Optimization for Vehicle-Home-Grid Integration including Household Load Prediction and Battery Degradation

### v2

@@ III RESULTS AND DISCUSSION
    Vehicle-home-grid integration: With objective function in (1), this scenario aims to minimize the user energy costs and battery degradation by controlling the bidirectional energy flows among the vehicle, household, and electric grid.
@@
    $\min\sum_{t}(E^{G2V}_{t}+E^{G2H}_{t})\cdot p_{t}+BC_{t}+s_{t}.$ (23)
III-A Cost and Battery Degradation Analysis
With the price ratio $\gamma=1$, meaning the price for purchasing and selling energy is identical, Table I summarizes the user costs for the two scenarios. Here, the user’s final cost ($FC$) is calculated as the sum of the energy cost ($EC$) and battery degradation cost ($BC$). $BD^{cyc}$ is the sum of cycle aging under high temperature, low temperature, and high temperature with low SoC. $E_{batt}$ represents the total energy that flows in and out of the EV battery.
TABLE I: User costs and battery degradation for scenarios A and B.
$FC$ $EC$ $BC$ $BD$ $BD^{cal}$ $BD^{cyc}$ $E_{batt}$
@@ III-A Cost and Battery Degradation Analysis
It can be observed that scenario A yields favorable performance: the negative energy cost ($EC$) indicates profit generated by selling electricity back to the grid via V2G. Correspondingly, the EV battery has degraded 5.42%. Overall, scenario A achieve an annual user profit ($FC$) of €1070.21.
@@
By comparing scenarios A and B, it can be concluded that the vehicle-home-grid integration provides a substantial economic advantage of €3046.81 annually for the user. Specifically, scenario A degrades the battery by 1.96% more but reduces the energy costs by €3288.50 compared to scenario B.
III-B Sensitivity Analysis of Price Ratio $\gamma$
@@
Figure 5: $FC$ for scenario A and B varying the price ratio $\gamma$.
As expected, in scenario B, the curve remains constant due to unidirectional charging, which does not involve V2G and thus is independent of the price ratio $\gamma$. In scenario A, for $\gamma=1$, $FC$ corresponds to the value shown in Table I. As $\gamma$ decreases, $FC$ increases because the reduced price difference between buying and selling makes V2G less profitable. For $\gamma=0.75$, $FC=0$. This indicates that using the EV for vehicle-home-grid integration eliminates EV-related costs, corresponding to an economic gain of €1976.60 that the user would have spent on unidirectional charging. However, for $\gamma<0.75$, $FC$ becomes positive, meaning that while the economic gain persists, no further profit is generated. When $\gamma=0$, V2G is no longer performed as it offers no benefit; however, V2H continues, contributing to self-consumption. This results in an economic gain (corresponding to a saving) of €425.48 compared to scenario B.
@@
III-C Impact of Battery Capacity & Household Load Variations
@@
When the consumption is $HL$ (solid lines), the curves for different $E_{b}$ values exhibit a clear increasing slope with $\gamma$, due to more effective V2G utilization. At low $\gamma$ values (less than 0.2), the economic gain decreases with larger batteries because any of considered battery capacities is sufficient to cover this $HL$, whereas larger batteries lead to increased battery replace costs, i.e., $C_{rep}$ in (4). Conversely, at higher $\gamma$ values, larger batteries result in greater economic gains due to enhanced V2G opportunities.
@@
With increased household consumption ($4HL$ and $8HL$), the economic gain curves become flatter as $\gamma$ increases, indicating reduced V2G opportunities due to the higher household energy demand. Specifically, for $4HL$ (dashed lines), a 41 kWh battery is clearly limited and yields the smallest economic gain among all tested battery capacities. Moreover, in the $4HL$ case, smaller batteries offer slightly higher gains when $\gamma$ is low, while the opposite trend is observed at higher $\gamma$, consistent with the behavior seen in the $HL$ case. In the case of $8HL$ (dash-dot lines), the very high household load severely limits the availability of energy for grid export, effectively suppressing V2G utilization. Unlike the $HL$ and $4HL$ cases, a larger battery consistently yields higher economic gain across the entire $\gamma$ range. This is because larger batteries provide the capacity needed to handle the high household demand more flexibly, even if V2G is minimally utilized.
@@
Additionally, for $\gamma=1$ and the same $E_{b}$, the economic gain remains constant across different household loads. This occurs because, with $\gamma=1$, there is no financial penalty on energy sales, and the user’s benefit depends solely on battery capacity. As a result, while a higher household load increases $FC$ in scenarios A and B, their economic difference remains unchanged for the same battery capacity.
@@
Overall, results in Fig. 6 confirm that vehicle-home-grid integration consistently provides economic benefits across all tested cases. Specifically, regardless of the price ratios, battery capacities, and household loads, even in the least favorable conditions, the user achieves financial advantages compared to the unidirectional charging in scenario B.
IV CONCLUSION
@@
This paper proposed a practical, real-time implementable optimization algorithm to manage the energy flows of vehicle-home-grid integration. The contributions first arise from the explicit consideration of nonlinear battery cycle and calendar aging, uncertainties in future household energy consumption, and a set of constraints in vehicle usage and battery dynamics. By utilizing a hybrid LSTM neural network to predict household loads and a detailed battery model to describe battery behavior, this algorithm results in minimized user energy and battery costs. This responds to a maximum annual economic gain of €3046.81 for a single user over the benchmark scenario with unidirectional smart charging.
@@
The second major contribution is the conducted systematic analysis of various price ratios, battery capacities, and household loads within the optimization problem. We have found that

### v3

@@ 3 RESULTS AND DISCUSSION
    Proposed vehicle-home-grid integration: With objective function in (1), this scenario aims to minimize the user energy costs and battery degradation by controlling the bidirectional energy flows among the vehicle, household, and electric grid.
@@
    $\min\sum_{t}(E^{G2V}_{t}+E^{G2H}_{t})\cdot p_{t}+BC_{t}.$ (23)
3.1 Cost and Battery Degradation Analysis
With the price ratio $\gamma=1$, meaning the price for purchasing and selling energy is identical, Table 1 summarizes the user costs for the two scenarios. Here, the user’s final cost ($FC$) is calculated as the sum of the energy cost ($EC$) and battery degradation cost ($BC$). $BD^{cyc}$ is the sum of cycle aging under high temperature, low temperature, and high temperature with low SoC. $E_{batt}$ represents the total energy that flows in and out of the EV battery.
Table 1: User costs and battery degradation for scenarios A and B.
FCFC ECEC BCBC BDBD $BD^{cal}$ $BD^{cyc}$ $E_{batt}$
@@ 3.1 Cost and Battery Degradation Analysis
It can be observed that scenario A yields favorable performance: the negative energy cost ($EC$) indicates profit generated by selling electricity back to the grid via V2G. Overall, scenario A achieve an annual user profit of €1070.21.
@@
By comparing scenarios A and B, it can be concluded that vehicle-home-grid integration provides an annual economic advantage of €3046.81 for the user. Specifically, scenario A degrades the battery by 1.96% more but reduces the energy costs by €3288.50 compared to scenario B.
3.2 Sensitivity Analysis of Price Ratio $\gamma$
@@
Figure 4: FCFC for scenario A and B varying price ratio $\gamma$.
As expected, in scenario B, the curve remains constant due to unidirectional smart charging, which does not involve V2G and thus is independent of the price ratio $\gamma$. In scenario A, for $\gamma=1$, $FC$ corresponds to the value shown in Table I. As $\gamma$ decreases, $FC$ increases because the reduced price difference between buying and selling makes V2G less profitable. For $\gamma=0.75$, $FC=0$. This indicates that using the EV for vehicle-home-grid integration eliminates EV-related costs, corresponding to an economic gain of €1976.60 that the user would have spent on unidirectional smart charging. However, for $\gamma<0.75$, $FC$ becomes positive, meaning that while the economic gain persists, no further profit is generated. When $\gamma=0$, V2G is no longer performed as it offers no benefit; however, V2H continues, contributing to self-consumption. This results in an economic gain (corresponding to a saving) of €425.48 compared to scenario B.
@@
3.3 Impact of Battery Size & Household Load Variations
@@
When the consumption is $HL$ (solid lines), the curves for different $E_{b}$ values exhibit a clear increasing slope with $\gamma$, due to more effective V2G utilization. At low $\gamma$ values (less than 0.2), the economic gain decreases with larger batteries because any of considered battery capacities is sufficient to cover this $HL$, whereas larger batteries lead to increased battery replace costs, i.e., $C_{rep}$ in (4). Conversely, at higher $\gamma$ values, larger batteries result in greater economic gains due to enhanced V2G.
@@
With increased household consumption ($4HL$ and $8HL$), the economic gain curves become flatter as $\gamma$ increases, indicating reduced V2G opportunities due to the higher household demand. Specifically, for $4HL$ (dashed lines), a 41 kWh battery is clearly limited, giving the lowest economic gain among all tested battery sizes. Moreover, in the $4HL$ case, smaller batteries offer slightly higher gains when $\gamma$ is low, while the opposite occurs at higher $\gamma$, consistent with the behavior seen in the $HL$ case. For $8HL$ (dash-dot lines), the very high household load severely limits the availability of energy for grid export, suppressing V2G utilization. Unlike the $HL$ and $4HL$ cases, a larger battery consistently yields higher economic gain across the entire $\gamma$ range. This is because larger batteries provide the capacity needed to handle the high household demand more flexibly, even if V2G is minimally utilized.
@@
Additionally, for $\gamma=1$ and the same $E_{b}$, the economic gain remains constant across household loads. This occurs because, with $\gamma=1$, there is no financial penalty on energy sales, and the user’s benefit depends solely on battery capacity. As a result, while a higher household load increases $FC$ in scenarios A and B, their economic difference remains unchanged for the same battery size.
@@
Overall, results in Fig. 5 confirm that vehicle-home-grid integration consistently provides economic benefits in all tested cases. Regardless of price ratios, battery capacities, and household loads, even in the least favorable conditions, the user achieves financial advantages compared to the unidirectional smart charging scenario.
4 CONCLUSION
@@
This paper proposed a real-time optimization algorithm for managing energy flows in vehicle–home–grid integration. The contributions first arise from the explicit consideration of nonlinear battery cycle and calendar aging, uncertainties in future household load, and constraints on vehicle usage and battery dynamics. By utilizing a hybrid LSTM neural network to predict household loads and a detailed battery model, this algorithm results in minimized user energy and battery costs, yielding an annual economic gain of up to €3046.81 for a single user over the benchmark scenario with unidirectional smart charging.
@@
The second major contribution is a systematic analysis of various price ratios, battery capacities, and household loads within the optimization problem. We have found that